A baseball is thrown from a horizontal plane following a parabolic path with an initial velocity of 100 m/s at an angle of 30° above the horizontal. How far from the throwing point will the ball attain its original level (maximum Range)?
1
Expert's answer
2020-11-16T07:45:20-0500

Let's first find the time that the baseball takes to reach the maximum height from the kinematic equation:


"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."


Then, we can find the total flight time of the baseball:


"t=2t_{rise}=\dfrac{2v_0sin\theta}{g}."


Finally, we can find the maximum range of the baseball from the kinematic equation:


"R=v_0tcos\theta=v_0cos\theta\dfrac{2v_0sin\theta}{g}=\dfrac{v_0^2sin2\theta}{g},""R=\dfrac{(100\ \dfrac{m}{s})^2\cdot sin2\cdot 30^{\circ}}{9.8\ \dfrac{m}{s^2}}=884\ m."

Answer:

"R=884\ m."